OpenStudy (okdutchman7):

Please help. WILL MEDAL! A bus goes d miles in h hours and gets there 4 hours late. At what speed should the bus have been driving to have arrived on time? @Loser66 @surjithayer

OpenStudy (okdutchman7):

@Atrineas @Awolflover1 @CandyCove @dtan5457 @EclipsedStar

OpenStudy (okdutchman7):

@FortyTheRapper @inkyvoyd @JFraser @jim_thompson5910 @Miracrown

OpenStudy (okdutchman7):

@OpenStudyIsSoAwesome @rebeccaxhawaii @robtobey @sleepyjess @SolomonZelman

OpenStudy (okdutchman7):

@truwhovian11 @wolf1728 @zepdrix

OpenStudy (okdutchman7):

Speed my teacher refers to it as rate.

OpenStudy (okdutchman7):

I'm confused.

OpenStudy (okdutchman7):

I just don't understand after \[Speed = \frac{ Distance }{ Time }\]

OpenStudy (okdutchman7):

Yes, I understand that.

OpenStudy (okdutchman7):

yes

OpenStudy (okdutchman7):

\[9\div3\]

OpenStudy (okdutchman7):

\[Speed = \frac{ Distance }{ Time }\]

OpenStudy (okdutchman7):

yes

OpenStudy (okdutchman7):

Not yet. :3

OpenStudy (okdutchman7):

@aaronq @agent0smith @Directrix @ganeshie8 @gottennis121

OpenStudy (okdutchman7):

@HannahC234 @Hero @jabez177 @Jadeishere @phi

OpenStudy (okdutchman7):

@Preetha @purple_pink @Qwertty123 @radar @ShadowLegendX

OpenStudy (okdutchman7):

@TheSmartOne @Vocaloid @surjithayer

OpenStudy (okdutchman7):

@satellite73

OpenStudy (agent0smith):

Old speed \[\Large v_{old} = \frac{ d }{ h }\] New speed \[\Large v_{new} =\frac{ d }{h-4 }\] You can solve for d then use substitution to get an equation for the new speed in terms of the old one.

OpenStudy (anonymous):

Refer to the Mathematica attachment.